Required information Problem 9-1A Short-term notes payable transactions and entries LO P1 [The following information appliesJournal entry worksheet 4 5 6 7 1 2 3 8 Replaced the April 20 account payable to Locust with a 90-day, 8%, $ 35 ,000 note payJournal entry worksheet 1 2 4 5 6 7 8 Borrowed $54,000 cash from NBR Bank by signing a 120-day, 10%, $54,000 note payable. NoJournal entry worksheet 1 2 5 6 7 8 Paid the amount due on the note to Locust at the maturity date Note: Enter debits beforeJournal entry worksheet 1 2 3 6 7 Paid the amount due on the note to NBR Bank at the maturity date. Note: Enter debits beforeJournal entry worksheet 1 2 3 4 5 7 8 Borrowed $27,000 cash from Fargo Bank by signing a 60-day, 8 %, $27,000 note payable. NJournal entry worksheet < 1 2 3 4 5 6 7 8 Recorded an adjusting entry for accrued interest on the note to Fargo Bank. Note: EJournal entry worksheet 2 1 3 4 5 6 7 8 Paid the amount due on the note to Fargo Bank at the maturity date. Note: Enter debit
Answer #1
1) Maturity date
locust NBR fargo
date of the note 19-May 8-Jul 28-Nov
term of note 90 120 60
maturity date 17-Aug 5-Nov 27-Jan
2) interest due at maturity
principal * Rate * time = interest
locust 35,000 * 8% * 90/360 = 700
NBR 54,000 * 10% * 120/360 = 1800
Fargo 27,000 * 8% * 60/360 = 360
3) Amount in adjusting entry
42,000*8%*33/360
198
Fargo Bank
principal * Rate * time = interest
interest to be acccrued in 2016 27,000 * 8% * 33/360 = 198
4) interest expense to be recorded in 2017
198
principal * Rate * time = interest
interest to recorded in 2018 27,000 * 8% * 27/360 = 162
Journal entries
Date Accounting titles & Explanations Debit Credit
2016
20-Apr inventory 35,500
Accounts payable 35,500
19-May Accounts payable 35,500
cash 500
notes payable 35,000
8-Jul Cash 54,000
notes payable 54,000
17-Aug notes payable 35,000
interest expense 700
cash 35,700
5-Nov notes payable 54,000
interest expense 1,800
cash 55,800
28-Nov Cash 27,000
notes payable 27,000
31-Dec interest expense 198
interest payable 198
2017
27-Jan notes payable 27,000
interest payable 198
interest expense 162
cash 27,360